# Learning to combine top-down context and feed-forward representations under ambiguity with apical and basal dendrites

### v2

@@ Abstract
One of the hallmark features of neocortical anatomy is the presence of extensive top-down projections into primary sensory areas, with many impinging on the distal apical dendrites of pyramidal neurons. While it is known that they exert a modulatory effect, altering the gain of responses, their functional role remains an active area of research. It is hypothesized that these top-down projections carry contextual information that can help animals to resolve ambiguities in sensory data. One proposed mechanism of contextual integration is a non-linear integration of distinct input streams at apical and basal dendrites of pyramidal neurons. Computationally, however, it is yet to be demonstrated how such an architecture could leverage distinct compartments for flexible contextual integration and sensory processing when both sensory and context signals can be unreliable. Here, we implement an augmented deep neural network with distinct apical and basal compartments that integrates a) contextual information from top-down projections to apical compartments, and b) sensory representations driven by bottom-up projections to basal compartments. In addition, we develop a new multi-scenario contextual integration task using a generative image modeling approach. In addition to generalizing previous contextual integration tasks, it better captures the diversity of scenarios where neither contextual nor sensory information are fully reliable. To solve this task, this model successfully learns to select among integration strategies. Specifically, when input stimuli and contextual information are contradictory, the performance of deep neural networks augmented with our "apical prior" exceeds that of single-compartment networks with otherwise equivalent architecture. We further show that our model can learn to integrate contextual information across time by using a task with input sequences and a gated recurrent unit layer to generate contextual signals. Using layerwise relevance propagation, we extract the relevance of individual neurons to the model’s predictions, revealing that a sparse subset of neurons encoding features of the context-relevant categories receive the largest magnitude of top-down signals when context and sensory input are aligned. We further show that this sparse gain modulation is necessary for best performance on the task. Altogether, this suggests that the "apical prior" and the biophysically-inspired non-linear integration rule could be key components necessary for handling the ambiguities that animals encounter in the diverse contexts of the real world.
@@ 1 Introduction
Accurate perception relies on appropriate integration of context, as sensory signals contain incomplete or ambiguous information (Rao and Ballard, 1999; Mumford, 1992). Contextual information can be derived from a number of sources, including the spatial and temporal domains (Eichenbaum, 2017), task demands (Gilbert and Li, 2013), as well as other sensory modalities.
Several key neuroanatomical and cellular features could support the computations associated with context integration. In the mammalian brain, context has been proposed to be conveyed by top-down feedback pathways which are abundant in sensory regions of the neocortex (Felleman and Van Essen, 1991; Markov et al., 2014; Harris et al., 2019). Perturbations of top-down pathways from higher order areas have been associated with delayed object recognition (Kar and DiCarlo, 2021) as well as disrupted stimulus response curves in lower-order regions (Wang et al., 2007; Nassi et al., 2013), which may reflect deficits in contextual processing. Several feed-forward computational models of perception have implemented top-down pathways to enable object recognition of occluded images (George et al., 2017; Spoerer et al., 2017) as well as context-invariant perception (Naumann et al., 2022), and context-dependent gating for multi-task learning Masse et al. (2018).
In the neocortex, pyramidal neurons are thought to integrate feedforward and top-down, feedback information at their basal and apical dendrites, respectively (Larkum et al., 2007; Spruston, 2008). Notably, feedback has been proposed to modulate rather than drive neuronal activities (Sherman and Guillery, 1998). This idea is supported by the unique morphological and physiological properties of pyramidal neurons, as activity in apical dendrites in vivo is generally not sufficient to drive somatic output responses alone (Stuart and Spruston, 1998; Larkum et al., 1999; Larkum, 2013; Jarvis et al., 2018). However, activation of the apical compartment has been shown to act as a gain modulator, amplifying concurrent basal activity (Larkum et al., 2004), serving as a potential stream that could carry contextual information. While the physiological properties of pyramidal neurons have been described extensively (see Spruston (2008) for review), there is no consensus on the exact mechanism by which top-down signals from higher-order regions modulate somatic firing rates and update sensory representations.
Here, we propose a deep neural network architecture in which neurons are augmented specifically with apical and basal compartments and whose computation respects known biophysical properties of pyramidal neurons. We design a training paradigm such that, according to the context, the model updates its higher-level representations of sensory inputs, via the interaction of basal and apical compartments, in which the latter carries top-down contextual information.
@@
We show that the model learns to use apical modulation, driven by contextual inputs, to resolve sensory ambiguities. We find that a neuron-wise (element-wise) dendritic integration rule informed by physiological data enables strong performance in both settings of the task.
To gain insights into the learned mechanisms of top-down modulation, we applied a standard method from the explainable AI literature ((Bach et al., 2015)) to identify the relative contributions of individual neurons and their apical compartments to network function. We identified a subset of neurons that are highly relevant for the matching category, and selectivelyamplified by top-down signals arriving at their apical dendrites.
We propose that, in principle, 1) apical gain modulation of somatic activity and contextual integration are compatible, and is a simple yet efficient way to solve new tasks such as resolving perceptual ambiguity without altering what the model has previously learned, as it does not introduce or modify any synapses in the feed-forward pathway, and 2) that the apical gain modulatory signals and the representations of contextual information arriving at the aforementioned apical compartments can be decoupled. In a highly distributed system with local learning constraints such as the brain, the representation of context can be learned independently from the mapping of contextual signals to local populations to solve specific perceptual tasks.
@@ 2.1 Functional model of context integration in apical dendrites
To model how apical dendrites integrate context to refine perceptual representations, we first developed a multi-scenario contextual integration task where correct classification of ambiguous characters requires appropriate integration of contextual cues (Fig. 1a). To create the task, we trained a generative model on handwritten digits and letters (MNIST and EMNIST, respectively) to construct a dataset with highly ambiguous characters. We used the generative model to synthesize images that are a mixture of two classes (see Methods). The ambiguity of the synthesized images was empirically quantified as the proportion of the confusion of a linear classifier trained on unambiguous images (fig. 1b). Inspired by anatomical features of pyramidal neurons, we next developed a model that implements distinct dendritic compartments, a basal compartment for integrating sensory information and an apical compartment for integrating contextual information (Fig. 1c). In the model, a population of $n$ pyramidal neurons has two sources of inputs, represented as vectors: a set of basal inputs, $\mathbf{b}=[b_{1},\ldots,b_{n}]$, representing the sensory stream, and a set of apical inputs, $\mathbf{a}=[a_{1},\ldots,a_{n}]$, representing the top-down stream. Input images, $\mathbf{x}$, are encoded by a pretrained network $f$, with parameters $\Theta_{b}$, where $f$ is a convolutional neural network reflecting initial stages of sensory processing (see Methods for more details on how the encoder is pretrained). The output of $f(x)$ is the basal vector $\mathbf{b}$.
@@
where $\odot$ is an neuron-wise multiplication and $\sigma$ is the rectified linear unit (ReLU) activation function implementing basal and apical non-linearities, respectively. As can be seen from equation 1, the impact of the apical input vector, $\mathbf{a}$, on the firing rate, $\mathbf{h}$, is determined by $\sigma(\mathbf{a})+1$. As such, the apical activity serves as a thresholded neuron-wise gain modulator of the basal activity, in-line with previous experimental reports of a multiplicative role for apical inputs (Waters et al., 2003). Under these constraints, apical inputs affect neuronal activity only when they exceed their activation threshold (when $a>0$), analogous to observations that apical inputs influence somatic spiking only if they evoke a dendritic calcium spike (Larkum et al., 1999). To monitor the ability of the model to solve the perceptual task, the output of the model is read-out from a linear transformation of the pyramidal neurons firing rates:
@@
where $\mu$ represents a higher-order neuronal population ($\mathbf{U}$ is a linear transformation to the latent space of the encoder $f$). To obtain the apical inputs $\mathbf{a}$, the context signal, $\mathbf{c}$, is concatenated with $\mu$, and mapped onto the apical compartment using a multilayer perceptron (MLP), $g(\cdot)$, as follows:
@@
As a population, these pyramidal neurons integrate sensory input character representations onto their basal dendritic compartment on one hand, and contextual information onto their apical dendritic compartment on the other hand.
For ambiguous characters, which are a mixture of two categories, a readout of the representations of the model will yield two plausible interpretations, and contextual information is required to resolve the ambiguity. We consider the category which is compatible with the contextual input to be the target category. Conversely, a category which is a plausible interpretation of the input but not compatible with the contextual information is referred to as the contradictory category. We refer to all other categories as irrelevant categories.
To test whether the model can learn apical modulations that are helpful for resolving ambiguous stimuli, we trained the synaptic weights $\Theta_{a}$ with the objective of modulating basal representations to match the target $\mu$ (representation after linear transform). We optimize this objective simultaneously for a set of 5 input-context scenarios (unambiguous or ambiguous input; helpful, contradictory or irrelevant context (Fig. 1e) using a common gradient descent optimization algorithm (Adam, see pseudocode 1, and 5). The labels and corresponding descriptions we use for each scenario are provided explicitly below:
@@
Figure 1: Functional model of context integration in apical dendrites. a, in our experimental paradigm, ambiguous characters (in yellow, top panels) are resolved by combining context signals (bottom panels). b, examples of unambiguous pairs and resulting ambiguous characters generated using EMNIST (left) and MNIST (right). c, model of pyramidal neurons with distinct dendritic compartments trained to solve a multi-scenario contextual integration task with ambiguity. Before top-down modulation, two predictions have equal probability. After top-down modulation, context contributes to favoring one unambiguous representation. d, implementation of dendritic integration in artificial neurons with apical and basal compartments. $f$ represents the feedforward mapping onto the basal compartment parameterized by the pretrained backbone, and $g$ the contextual top-down mapping onto the apical compartment, parameterized by its own neural network. The combination of the outputs $b$ and $a$ of $f$ and $g$, respectively, is applied neuron-wise via the multiplicative interaction $\sigma(b)*(1+\sigma(a))$. e, distinct combinations of input and context used to train the top-down network $g$. f, possible combinations of input and context used during training and analysis.
@@ 2.2 Resolving ambiguity in the oracle context
To first test if our model can solve our multi-scenario task under contextual input represented as a one-hot vector (oracle context), we trained $g$ on the ambiguous MNIST/EMNIST task with a loss function that includes a loss term for each scenario We assessed performance of the model in each scenario as the accuracy of the readout on the test set, and found significant impact of top-down modulation. (one-way ANOVA; $F_{4}$ = 10076.598; p $<$ 0.0001; n = 313 independent samples). Specifically, test set accuracy on ambiguous digits was $46.3\pm 0.3\%$ without top-down modulation, and became significantly higher with top-down modulation ($98.5\pm 0.3\%$; Tukey’s test; p $<$ 0.0001) if context was matching, but did not significantly alter performance if context was irrelevant ($43.920\pm 8.160$; Tukey’s test; p = 0.4651; n = 313 independent replicates; Fig. 2a; Table 2). Importantly, the model did not learn to rely solely on top-down context, since performance remained high when the model was given contradictory context (Fig. 2a; 2). To visualize the representations of the model on a population level, we projected unmodulated and modulated outputs $\mathbf{\mu}$ onto a two-dimensional manifold using t-SNE (Maaten (2008); Fig. 2b). Ambiguous images whose representations exhibited high overlap between class-pairs were effectively disentangled after integrating matching contextual inputs (paired t-test, t99 = 6.0293, p $<$ 0.0001 for silhouette scores between ambiguous images before and after integrating top-down inputs; Fig. 2c). These results show that our model learned to use contextual inputs provided to apical dendrites to modulate basal activity and resolve ambiguity.
Figure 2: Our task is best solved when top-down signals are integrated correctly. a, readout test set accuracy before and after combining top-down signals over training epochs for all input-context combinations. Dashed line corresponds to chance level. b, low dimensional projection (t-SNE) of $h$ for unambiguous inputs (top), ambiguous inputs before (bottom left), and after (bottom right) top-down modulation. c, silhouette scores for $h$, for each class pair and for unambiguous inputs (left), ambiguous inputs before (center), and after (right) top-down modulation. d, cosine distance between $h$ in unambiguous conditions and basal, apical compartments and firing rate in each scenario.
@@
Next, we evaluated the importance of having topdown apical compartments to handle contextual inputs compared with only having basal compartments (akin to point-wise spiking neurons, (Table. 4). We found that, while there was not a large difference in the other 4 scenarios, there was an average increase in performance of 1.9% (MNIST) in the unambig.Contra scenario for the models which had apical compartments compared to those which only had basal compartments.
Next, focusing our analysis on MNIST, we evaluated the importance of the integration rule which combines basal and apical activities4, 3a).
@@ 2.3 Single neuron mechanisms of top-down modulation
We next sought to understand the computational role of top-down modulation of pyramidal apical dendrites in solving this contextual integration task. To this end, we used a method from the explainable AI literature to rank and identify the neurons most relevant for solving the task. We then compared their activations across the various context-input scenarios. First, to define the most relevant neurons, we employed layer-wise relevance propagation (LRP; (Bach et al., 2015); supplementary Fig. 1a), a common method in explainable AI literature to explain, at each layer, the input’s contribution to the model output. Each neuron receives a score for its contribution to the model output for each image (Supplementary Fig. 1b). We then defined the subset of neurons most relevant for identifying a given class, $S_{C}$, as the minimum set of neurons required to account for 95% of total neuronal relevance when processing images from that class (Fig.  1c,d, see Methods). Sets of relevant neurons defined in this way exhibit low overlap between classes (Supplementary Fig. 1b-d, see Methods). We observed similar representational separability between classes for other scenarios and given appropriate top-down inputs (Supplementary Fig. 1e). Additionally, these sets of relevant neurons are typically sparse and represent $\sim$15of the total neuronal population (Supplementary Fig. 1b,c).
@@
It is noteworthy that our model is not explicitly provided with the ambiguous nature of input images, and therefore learns to extract this information to solve the task. In addition, the model learns to ignore top-down signals when the image is unambiguous and context is contradictory or irrelevant. Given that the apical compartment must combine contextual information and the contextually naive $\mu$ representation differentially based on input ambiguity, we hypothesized that solving the task effectively required the gain modulation of a subset of neurons encoding features specific to the relevant class. To assess this possibility, we registered the amplitude of apical signals arriving at LRP subsets pertaining to class-specific neurons and class-pair overlapping neurons in each scenario (Fig. 4a-f). When comparing all input/context scenarios, we found that apical signals were highest when input images were ambiguous and context informative (1-ANOVA, F416=237.024, p $<$ 0.0001, for the main effect of input/context scenario; Fig. 4b). In the ambig.Match scenario, neurons most relevant for the target class were associated with the highest amplitude in the apical compartment (1-ANOVA, F417=17.9445, p $<$ 0.0001, for the main effect of neuron class; Fig. 4b). We hypothesized that target class specific neurons are also most modulated by top-down context, and that modulating these neurons is necessary to solve the perceptual ambiguity classification task.(Fig. 4c).
@@
To evaluate this hypothesis, we applied a mask on selected neurons post-training, and assessed the model’s ability to solve the task. Specifically, we compare test set accuracy when inputs are ambiguous and context is relevant, under the condition where neurons (as identified with LRP) encoding features specific to the relevant class (which we refer to as context-relevant neurons) are masked out, compared to the condition where random neurons are masked. We also evaluate the effect of mask size and found that masking the context-relevant neurons specifically led to a steeper degradation of accuracy as we increase the number of masked neurons. (Fig. 4d). This highlights the key role of these neurons’ apical compartments in integrating relevant contextual signals.
Figure 4: Top-down has information about ambiguity, and in cases where inputs are ambiguous and context is aligned, resolves ambiguity by amplifying context-relevant neurons.. a, using layer-wise relevance propagation, we identify neurons most relevant for either of two classes (red and blue), or overlap of both classes (magenta). b, amplitude of the apical compartment (normalized over all neurons) for each input/context condition. Three distinct classes of neurons are compared: neurons most relevant for unambiguous class A (blue), B (red), or neurons most relevant for both A and B classes (magenta). c, we hypothesize that context-relevant neurons receive the highest amplitude of top-down signals when processing ambiguous images, especially given matching context.

### v3

@@ Abstract
One of the hallmark features of neocortical anatomy is the presence of extensive top-down projections into primary sensory areas, with many impinging on the distal apical dendrites of pyramidal neurons. While it is known that they exert a modulatory effect, altering the gain of responses, their functional role remains an active area of research. It is hypothesized that one of the roles of these top-down projections is to carry contextual information that can help animals to resolve ambiguities in sensory data. One proposed mechanism of contextual integration is a combination of input streams at distinct (separate integration zone) apical and basal dendrites of pyramidal neurons. Computationally, however, it is yet to be demonstrated how such an architecture could leverage distinct compartments for flexible contextual integration and sensory processing when both sensory and context signals can be unreliable. Here, we implement an augmented deep neural network with distinct apical and basal compartments that integrates a) contextual information from top-down projections to apical compartments, and b) sensory representations driven by bottom-up projections to basal compartments. In addition, we develop a new multi-scenario contextual integration task using a generative image modeling approach. In addition to generalizing previous contextual integration tasks, it better captures the diversity of scenarios where neither contextual nor sensory information are fully reliable. To solve this task, this model successfully learns to select among integration strategies. Specifically, when input stimuli and contextual information are contradictory, the performance of deep neural networks augmented with our "apical prior" exceeds that of single-compartment networks with otherwise equivalent architecture. We further show that our model can learn to integrate contextual information across time by using a task with input sequences and a gated recurrent unit layer to generate contextual signals. Using layerwise relevance propagation, we extract the relevance of individual neurons to the model’s predictions, revealing that a sparse subset of neurons encoding features of the context-relevant categories receive the largest magnitude of top-down signals when context and sensory input are aligned. We further show that this sparse gain modulation is necessary for best performance on the task. Altogether, this suggests that the "apical prior" and the biophysically-inspired non-linear integration rule could be key components necessary for handling the ambiguities that animals encounter in the diverse contexts of the real world.
@@ 1 Introduction
Accurate perception relies on appropriate integration of context, as sensory signals contain incomplete or ambiguous information (Mumford, 1992; Rao and Ballard, 1999). Contextual information can be derived from a number of sources, including the spatial and temporal domains (Eichenbaum, 2017), task demands (Gilbert and Li, 2013), as well as other sensory modalities.
Several key neuroanatomical and cellular features could support the computations associated with context integration. In the mammalian brain, context has been proposed to be conveyed by top-down feedback pathways which are abundant in sensory regions of the neocortex (Felleman and Van Essen, 1991; Markov et al., 2014; Harris et al., 2019). Perturbations of top-down pathways from higher order areas have been associated with delayed object recognition (Kar and DiCarlo, 2021) as well as disrupted stimulus response curves in lower-order regions (Wang et al., 2007; Nassi et al., 2013), which may reflect deficits in contextual processing. Several feed-forward computational models of perception have implemented top-down pathways to enable object recognition of occluded images (George et al., 2017; Spoerer et al., 2017) as well as context-invariant perception (Naumann et al., 2022), and context-dependent gating for multi-task learning (Masse et al., 2018).
In the neocortex, pyramidal neurons are thought to integrate feedforward and top-down, feedback information at their basal and apical dendrites, respectively (Larkum et al., 2007; Spruston, 2008). Notably, feedback has been proposed to modulate rather than drive neuronal activities (Sherman and Guillery, 1998). This idea is supported by the unique morphological and physiological properties of pyramidal neurons, as activity in apical dendrites in vivo is generally not sufficient to drive somatic output responses alone (Stuart and Spruston, 1998; Larkum et al., 1999; Larkum, 2013; Jarvis et al., 2018). However, activation of the apical compartment has been shown to act as a gain modulator, amplifying concurrent basal activity (Larkum et al., 2004), serving as a potential mechanism for contextual information. While the physiological properties of pyramidal neurons have been described extensively (see Spruston (2008) for review), there is no consensus on the exact mechanism by which top-down signals from higher-order regions modulate somatic firing rates and update sensory representations.
Here, we propose a deep neural network architecture in which neurons are augmented specifically with apical and basal compartments and whose computation respects known biophysical properties of pyramidal neurons. We design a training paradigm such that the model updates its representations of sensory inputs arriving at the basal compartment according to the top-down contextual information arriving at the apical compartment.
@@
We show that the model learns to use apical modulation, driven by contextual inputs, to resolve sensory ambiguities. To gain insights into the learned mechanisms of top-down modulation, we applied a standard method from the explainable AI literature ((Bach et al., 2015)) to identify the relative contributions of individual neurons and their apical compartments to network function. We identified a subset of neurons that are highly relevant for the contextually-relevant category, and selectively amplified by top-down signals on apical dendrites.
Our results show that apical gain modulation of somatic activity by contextual information is a simple yet efficient way to solve new tasks, such as resolving perceptual ambiguity. Moreover, the approach is flexible, because it is achieved without altering what the model has previously learned, as it does not introduce or modify any synapses in the feed-forward pathway.
@@ 2.1 Functional model of context integration in apical dendrites
To model how apical dendrites integrate context to refine perceptual representations, we first developed a multi-scenario contextual integration task where correct classification of ambiguous characters requires appropriate integration of contextual cues (Fig. 1a). To create the task, we trained a generative model on handwritten digits and letters (MNIST and EMNIST, respectively) to construct a dataset with highly ambiguous characters. We used the generative model to synthesize images that are a mixture of two classes (see Methods). The ambiguity of the synthesized images was empirically quantified via softmax on classifier predictions, where we select images near 50% for two classes (fig. 1b). Inspired by anatomical features of pyramidal neurons, we next developed a model that implements distinct dendritic compartments, a basal compartment for integrating sensory information and an apical compartment for integrating contextual information (Fig. 1c). In the model, a population of $n$ pyramidal neurons has two sources of inputs, represented as vectors: a set of basal inputs, $\mathbf{b}=[b_{1},\ldots,b_{n}]$, representing the sensory stream, and a set of apical inputs, $\mathbf{a}=[a_{1},\ldots,a_{n}]$, representing the top-down stream. Input images, $\mathbf{x}$, are encoded by a pre-trained network $f$, with parameters $\Theta_{b}$, where $f$ is a convolutional neural network reflecting initial stages of sensory processing (see Methods for more details on how the encoder is pre-trained). The output of $f(x)$ is the basal vector $\mathbf{b}$.
@@
where $\odot$ is an neuron-wise multiplication (Hadamard) and $\sigma$ is the rectified linear unit (ReLU) activation function implementing basal and apical non-linearities, respectively. As can be seen from equation 1, the impact of the apical input vector, $\mathbf{a}$, on the firing rate, $\mathbf{h}$, is determined by $\sigma(\mathbf{a})+1$. As such, the apical activity serves as a thresholded neuron-wise gain modulator of the basal activity, in-line with previous experimental reports of a multiplicative role for apical inputs (Waters et al., 2003), an architectural bias of neocortical networks we refer to here as the "apical prior." Under these constraints, apical inputs affect neuronal activity only when they exceed their activation threshold (when $a>0$), analogous to observations that apical inputs influence somatic spiking only if they evoke a dendritic calcium spike (Larkum et al., 1999). Thus, this neuron-wise gain modulation of basal representations by the apical compartment via the Hadamard product is an implementation of the "apical prior."
To monitor the ability of the model to solve the perceptual task, we read-out model predictions from a linear transformation of the pyramidal neurons firing rates:
@@
where $\mu$ represents a higher-order neuronal population and $\mathbf{U}$ is a linear transformation (to the latent space of the encoder $f$). To obtain the apical inputs $\mathbf{a}$, the context signal, $\mathbf{c}$, is concatenated with $\mu$, and mapped onto the apical compartment using a multi-layer perceptron (MLP), $g(\cdot)$, as follows:
@@
As a population, these pyramidal neurons integrate sensory inputs (representations of characters) onto their basal dendritic compartment on one hand, and contextual information onto their apical dendritic compartment on the other hand.
For ambiguous characters, which are a mixture of two categories, a readout of the representations of the model will yield two plausible interpretations, and contextual information is required to resolve the ambiguity. We consider the category which is compatible with the contextual input to be the target (match) category. Conversely, a category which is a plausible interpretation of the input but not compatible with the contextual information is referred to as the contradictory category. We refer to all other categories as irrelevant categories. In total, we consider 5 possible scenarios. The labels and corresponding descriptions we use for each scenario are provided explicitly below:
@@
To test whether the contextual integration model incorporating the "apical prior" can learn apical modulations that are helpful for resolving ambiguous stimuli, we trained the synaptic weights $\Theta_{a}$ with the objective of modulating basal representations to match the target $\mu$ (representation after linear transform). We optimize this objective simultaneously for the set of all input-context scenarios listed above (Fig. 1e and Eq.5) using a gradient descent optimization algorithm (Adam, see pseudocode 1).
@@
Figure 1: Functional model of context integration in apical dendrites. a, in our experimental paradigm, ambiguous characters (in yellow, top panels) are resolved by combining context signals (bottom panels). b, examples of unambiguous pairs and resulting ambiguous characters generated using EMNIST (left) and MNIST (right). c, model of pyramidal neurons with distinct dendritic compartments trained to solve a multi-scenario contextual integration task with ambiguity. Before top-down modulation, two predictions have equal probability. After top-down modulation, context contributes to favoring one unambiguous representation. d, implementation of dendritic integration in artificial neurons with apical and basal compartments. $f$ represents the feedforward mapping onto the basal compartment parameterized by the pre-trained backbone, and $g$ the contextual top-down mapping onto the apical compartment, parameterized by its own neural network. The combination of the outputs $b$ and $a$ of $f$ and $g$, respectively, is applied neuron-wise via the multiplicative interaction $\sigma(b)*(1+\sigma(a))$. e, distinct combinations of input and context used to train the top-down network $g$. f, possible combinations of input and context used during training and analysis.
@@ 2.2 Resolving ambiguity in the oracle context
To first test if our model can solve our multi-scenario task under contextual input represented as a one-hot vector (oracle context), we trained $g$ on the ambiguous MNIST/EMNIST task with a loss function that includes a loss term for each scenario We assessed performance of the model in each scenario as the accuracy of the readout on the test set, and found significant impact of top-down modulation. (one-way ANOVA; $F_{4}$ = 10076.598; p $<$ 0.0001; n = 313 independent samples). Specifically, test set accuracy on ambiguous digits was $46.3\pm 0.3\%$ without top-down modulation, and became significantly higher with top-down modulation ($98.5\pm 0.3\%$; Tukey’s test; p $<$ 0.0001) if context was matching, but did not significantly alter performance if context was irrelevant ($43.920\pm 8.160$; Tukey’s test; p = 0.4651; n = 313 independent replicates; Fig. 2a; Table 2). Importantly, the model did not learn to rely solely on top-down context, since performance remained high when the model was given contradictory context (Fig. 2a; 2). To visualize the representations of the model on a population level, we projected un-modulated and modulated outputs $\mathbf{\mu}$ onto a two-dimensional manifold using t-SNE (Maaten (2008); Fig. 2b). Ambiguous images whose representations exhibited high overlap between class-pairs were effectively disentangled after integrating matching contextual inputs (paired t-test, t99 = 6.0293, p $<$ 0.0001 for silhouette scores between ambiguous images before and after integrating top-down inputs; Fig. 2c). These results show that our model learned to use contextual inputs provided to apical dendrites to modulate basal activity and resolve ambiguity.
Figure 2: Our task is best solved when top-down signals are integrated correctly. a, readout test set accuracy without (black) and with (cyan) top-down signals for all input-context combinations. Dashed line corresponds to chance level. b, low dimensional projection (t-SNE) of the latent representation, $\mu$, for unambiguous inputs (top), ambiguous inputs before (bottom left), and after (bottom right) top-down modulation. c, silhouette scores for latent representation, $\mu$, for each class pair and for unambiguous inputs (left), ambiguous inputs before (center), and after (right) top-down modulation. d, cosine distance between unambiguous activation, $h$, and basal, apical compartments and firing rate in each scenario.
@@
Next, we evaluated the importance of having top-down apical compartments to handle contextual inputs compared with only having basal compartments (akin to point-wise spiking neurons, (Table. 4). We found that, while there was not a large difference in the other 4 scenarios, there was an average increase in performance of 1.9% (MNIST) in the unambig.Contra scenario for the models which had apical compartments compared to those which only had basal compartments.
Next, focusing our analysis on MNIST, we evaluated the importance of the integration rule which combines basal and apical activities (Fig. 3a, Appendix 1 Table. 4).
@@ 2.3 Single neuron mechanisms of top-down modulation
We next sought to understand the computational role of top-down modulation of pyramidal apical dendrites in solving this contextual integration task. To this end, we used a method from the explainable AI literature to rank and identify the neurons most relevant for solving the task. We then compared their activations across the various context-input scenarios. First, to define the most relevant neurons, we employed layer-wise relevance propagation (LRP; (Bach et al., 2015); Appendix Fig. 1a), a method in the explainable AI literature to reveal, at each layer, the input’s contribution to the model output. Each neuron receives a score for its contribution to the model output for each image (Appendix Fig. 1b). We then defined the subset of neurons most relevant for predicting a given class, e.g. class A or class B, as the minimum set of neurons required to account for 95% of total neuronal relevance when processing images from that class (Appendix Fig.  1c,d, see Methods). Similarly, we define subsets of neurons most relevant for predicting that an image is ambiguous between two classes A and B (ambiguous AB). Sets of relevant neurons defined in this way exhibit low overlap between classes (Appendix Fig. 1b-d, see Methods). We observed similar representational separability between classes for other scenarios (Appendix Fig. 1e). Additionally, these sets of relevant neurons are typically sparse and represent $\sim$15of the total neuronal population (Appendix Fig. 1b,c).
@@
It is noteworthy that our model is not explicitly provided with the ambiguous nature of the input images, and therefore learns to extract this information to solve the task. In addition, the model learns to ignore top-down signals when the image is unambiguous and context is contradictory or irrelevant. Given that the apical compartment must combine contextual information and the contextually naive $\mu$ representation differentially based on input ambiguity, we hypothesized that solving the task effectively required a high gain modulation specific to the subset of context-relevant neurons. To assess this possibility, we computed the average amplitude of apical signals arriving at LRP subsets pertaining to predicting each of two input classes (A,B) and ambiguous AB for each scenario of matching, irrelevant or contradictory context (Fig. 4a-f). When comparing all input/context scenarios, we found that apical signals were highest when input images were ambiguous and context informative (1-ANOVA, F416=237.024, p $<$ 0.0001, for the main effect of input/context scenario; Fig. 4b). In the ambig.Match scenario, neurons most relevant for the target class were associated with the highest amplitude in the apical compartment (1-ANOVA, F417=17.9445, p $<$ 0.0001, for the main effect of neuron class; Fig. 4b). Apical input onto neurons outside of these defined subpopulations ("other") remained low ($\sim$ 0.07-0.08) across all scenarios (data not shown), indicating there was no relationship between other neurons’ apical signal and the scenario.
We also computed the basal amplitude for the same subsets of neurons defined above and found that there was a higher average amplitude on the matching relative to the contradictory subpopulation for unambiguous stimuli. (Fig. 4c). Likewise, there was a higher basal amplitude on the ambiguous subpopulation relative to other subpopulations for ambiguous inputs. Basal input onto the "other" subpopulation was found to be significantly lower for ambig.match and unambig.contra scenarios, but similar in magnitude to the unambig/ambig.irrel scenarios ($<$0.02, Fig. 4c).
These observations are indicative of a mechanism whereby ambiguous bottom-up representations are biased towards top-down representations by specific apical inputs Fig. 4d, whereas unambiguous bottom-up representations are preserved due to non-specific top-down modulation Fig. 4e.
@@
To evaluate if specific top-down modulation is necessary for solving the ambig.match task, we applied a mask on the apical inputs to the corresponding subpopulations post-training, and assessed the model’s ability to solve the task (Appendix 2-Fig. 1). Specifically, we compare test set accuracy when inputs are ambiguous and context is relevant, under the condition where context-relevant apical inputs are masked out (activations set to 0), compared to the condition where random apical inputs are masked. We also evaluate the effect of mask size and found that masking the apical inputs to context-relevant neurons specifically led to a steeper degradation of accuracy as we increase the number of masked inputs. This highlights the key role of these neurons’ apical compartments in integrating relevant contextual signals.
Figure 4: Top-down has information about ambiguity, and in cases where inputs are ambiguous and context is aligned, resolves ambiguity by amplifying context-relevant neurons. a, using layer-wise relevance propagation, we identify neurons most relevant for either of two classes (red and blue), or for ambiguous inputs between the two classes (magenta). b, amplitude of the apical compartment (normalized over all neurons) for each input/context condition. Three distinct classes of neurons are compared: neurons most relevant for unambiguous class A (blue), B (red), or neurons most relevant for the ambiguous AB images (magenta). c, Normalized basal signal. left: in case of ambiguous stimuli, both plausible class-subpopulations are active along with the ambiguity detector subpopulation. right: in case of unambiguous stimuli, we observe a similar distribution, with the context-relevant neurons (class A) having higher average firing rate. d, a cartoon summary of the observed mechanism for the ambig.match-B case (ambiguous input, contextual input favouring class B), where class B neurons code for ’3’ (dark red) and class A neurons code for ’8’ (blue). Context-relevant neurons (dark red) receive the highest amplitude of top-down signals when processing ambiguous input. At basal dendrites, these sub-populations have low but non-zero firing rates, which allows them to be selectively amplified with top-down signals via the Hadamard integration rule, and thus dictate the output representation. e, for unambiguous input and contradictory context (unambig.contra), top-down input becomes non-specific and neurons are weakly amplified by the Hadamard rule. The output is then dictated by the neurons receiving the strongest basal input, in this case class B neurons (dark red). Output representations for unambiguous inputs are thus preserved when contextual inputs are added.